owl shift report, 08/18/99

Valeri Jejer (jejer@uvahea.phys.virginia.edu)
Wed, 18 Aug 1999 08:17:35 -0400 (EDT)

Shift Summary for owl shift, 18/Aug/99

On Shift: Hogan, Valeri.

Hours of beam available at all intensities: 5.6
Hours of beam available at nominal intensity: 1.0
Hours of E832 data (special runs not included): 5.6
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

Accelerator problems, kfrends.

========================================================================
Runs ended this shift:
None
(Run 14257 - in progress - BEAM_832)

Tapes brought to FCC this shift:
QKC(K,S)279-281, QKB098
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 datataking with a lot of beam aborts and low intensity (4e12) most
of the run.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14257 1 00:00 Start E832 run with kfrend after ktevana errors.
5 00:08 Live time = 0, DYC32 is busy -> kfrend.
53 00:57 Ktevana errors - unknown crates in stream 5 ->
kfrend.
01:03 MCR calls reducing intensity due to multiple aborts.
01:12 Mounted split tape KQSM12 on ktev4.
89 01:58 Missing FFFF word in crate 53 - kfrend.
02:25 MCR called saying they going to increase intensity,
beam aborts during the following spill. This was the
last try to increase the intensity.
02:10 CsI 226 has gain drift -4.8% vs database and stays
around through the rest of the shift.
02:47 CsI 1777 has Laser stab. RMS 33.8%, OK next spill.
02:49 Quench at injection, lost beam.
03:54 Beam is back.
05:35 Lost beam.
05:43 Beam is back.
05:55 CsI 1137 has Laser stab. RMS 12%, OK next spill.
06:01 TeV RF problems, estimate 20 min.
06:07 Beam is back.
========================================================================
Follow up Notes on Previous Shift

Monitor CsI channels 1137, 128 -- Both were OK.
========================================================================
Items to Follow up on Next Shift